A cell of emf 3v an internal resistance 0.5ohms passes current through an external resistor of 7o.calculate the potential drop across the cell


1
Expert's answer
2021-09-15T11:39:03-0400

Explanations & Calculations


  • The potential drop across the cell is exactly the same as that drops across the external resistance if the connecting wires & other stuff do not contain resistance in them.
  • Therefore, just by applying "\small v=ir" to that external one, the needed quantity can be evaluated.
  • Before that, the current flowing through the circuit should be calculated & which can be done by applying Kirchhoff's rule to the circuit. Therefore,

"\qquad\qquad
\begin{aligned}
\small i &=\small \frac{E}{R+r}=\frac{3V}{0.5\Omega+7\Omega}=0.4\,A
\end{aligned}"

  • Then simply,

"\qquad\qquad
\begin{aligned}
\small V&=\small iR\\
\small \bold{V _{drop}}&=\small \frac{ER}{(R+r)}=0.4\,A\times7\Omega=\bold{2.8V}
\end{aligned}"


  • This is also obtained by the relationship "\small V_{drop} = E\pm ir_{internal}", + during the cell being charged & - during the normal service.